what is the slope intercept form of the standard form 1/3 x + y = 8

Found 2 solutions by Cromlix, dkppathak:
Answer by Cromlix(4381)

1/3 x + y = 8
This must be put in the form;
y = mx + c
y = 1/3x + 8
Or if you multiply through
by 3
3y = x + 24
The gradient - m - 1/3
Y intercept - c - (0,8)
Hope this helps
:-)

Answer by dkppathak(439)

equation of line passing with intercept on y axis
Y=Mx+c where m is known as slope and c is intercept on y
1/3 x + y = 8
y=-1/3x+8
by comparison we can say
M=-1/3
and c=8
ANSWER slope =-1/3 and intercepts 8